suppose the odds against an event are 5:2
a}what are the odds in favor of the event?
b} what is the probabilty of the event?

Answer by Theo(13342)

if odds against are 5:2, then odds in favor are 2:5 and probability against is 5:7 and probability for is 2:7.

basically:

odds in favor equals probability for divided by probability against.

odds against equals probability against divided by probability for.

odds in favor equals reciprocal of odds against.

odds against equals reciprocal of odds in favor.

let's say you have 7 marbles.

5 of them are white and 2 of them are black.

probability of drawing a white marble = 5/7.

probability of drawing a black marble = 2/7

odds in favor of drawing a white marble = probability for divided by probability against = (5/7) / (2/7) = (5/7) * (7/2) = (5/2)

odds against drawing a white marble = probability against divided by probability for = (2/7) / (5/7) = (2/7) * (7/5) = (2/5).